If you use WYSIWYG and "Formated" instead of "Normal" your logs will look ... well... "normal" --jill

January 12-24 SDT Run

12-13 Jan 2007 Clouds, wind, Didn't open

13-14 Jan 2007 Ice, Fog. Didn't open

14-15 Jan 2007 Ice on enclosure. Clear, but didn't open

15-16 Jan 2007 Photometric, bad seeing. Data taken for Ryan program ( 20070116.hsel: 15/16 January 2007 log)

16-17 Jan 2007 Data taken for Trilling program ( 20070117.hsel: 16/17 January 2007 log)

17-18 Jan 2007 Cloudy, no useful data. (Note added 20-Feb by O Kuhn: the detailed log shows that some data were taken for Trilling, Ryan projects and tried one obs for Milne at end of night - check detailed log more carefully to determine usefulness of data). ( 20070118.hsel: 17/18 January 2007 log)

18-19 Jan 2007 Cloudy, no useful data

19-20 Jan 2007 Blowing snow, then enclosure problems, never opened.

20-21 Jan 2007 High humidity, ice on everything, teasing glimpses of starlight. never opened.

21-22 Jan 2007 Winter storm, never opened.

22-23 Jan 2007 Winter storm aftermath, fog and blowing snow, cryostat warmed up

23-24 Jan 2007 Photometric, seeing poor, high winds later, some observing!! 24-Jan-2007 image list

24-25 Jan 2007 Photometric, seeing awful, high winds forced closure shortly after midnight. 25-Jan-2007 image list

February 11-26 SDT Run

11-12 Feb 2007 Bad weather, never opened

12-13 Feb 2007 Bad weather, never opened

13-14 Feb 2007 Bad weather, never opened. Pinhole tests.

14-15 Feb 2007 Bad weather, never opened.

15-16 Feb 2007 Clear, but seeing poor: Milne SNe, M81, N5907 ( 20070216.hsel: 15/16 February 2007 log)

16-17 Feb 2007 Clear, seeing 1.2-1.6, SigOr, Cosmos_g, IC2574; lost end of night from tel. elevation encoders ( 20070217.hsel: 16/17 February 2007 log)

17-18 Feb 2007 Thin, seeing 1.2-1.5, SigOr, SNe, M81, M67, lost begining of night due to same problem; end lost due to weather ( 20070218.hsel: 17/18 February 2007 log)

18-19 Feb 2007 Open for nearly 4.5 hours; seeing 1.1-1.7, target in dwarf galaxy program nearly completed; M81, attempts at M67 and IC2574 ( 20070219.hsel: 18/19 February 2007 log)

19-20 Feb 2007 Bad weather, never opened.

20-21 Feb 2007 Clear night, little wind, seeing varied between 0.6 and 1.4" (typically was just a little worse than 1"). Seeing only <0.9" during last 1.5 hours of the night. Data for various programs were obtained. ( 20070221.hsel: 20/21 Feb 2007 log)

21-22 Feb 2007 Clear night at start. Great seeing at start but a few optical and guiding issues encountered. Seeing degraded significantly after midnight limiting observing options. Clouds start moving in around 0200. By 0300 it was impossible to observe. ( 20070222.hsel: 21/22 Feb 2007 log)

22-23 Feb 2007 Clear night at start and fairly windy because of an approaching front from California. Seeing generally around 1". Several major technical problems. 1) Telescope pointing was completely lost at start (many degrees off). As a result, the tracking was incorrect for the evening flats and stars are trailed according to the exposure length. It took about 1 hour to reset the telescope pointing. 2) The primary mirror got itself into a strange state with large plus-and-minus actuator forces along the outer edge of the mirror (astigmatism). This took almost 2 hours to sort through. Seeing worsen as front approaches from west. Third problem of the night: LBC control lost requiring reboot and restart of the instrument (nearly an hour spent trying to figure this one out). Mirror control problems come back. Not a very technically satisfying night. Lots of problems on many apparent fronts. Not too much accomplished scientifically except that KBO1 was completed. ( 20070223.hsel: 22/23 Feb 2007 log)

23-24 Feb 2007 Winter storm, never opened. Sky cleared, but humidity stayed too high throughout the night.

24-25 Feb 2007 Clear sky, but first quarter moon. Some early problems finding focus and the LBC crashing, but a productive night to finish the February 2007 SDT observing run. Thanks to the LBTO staff for all of the help in making this a successful run. ( 20070225.hsel: 24/25 Feb 2007 log)

25-26 Feb 2007 Some bonus SDT observing during engineering time. We are trying to finish up some M67 astrometric calibrations while the camera is still in the identical baffle configuration. Seeing was horrible ~3 arcsec with strong wind. ( 20070226.hsel: 25/26 Feb 2007 log)

26-27 Feb 2007 Some more bonus SDT observing during engineering time. This time we did finish the M67 astrometric calibrations. Seeing ~0.9 arcsec under photometric skies with moon. ( 20070227.hsel: 26/27 Feb 2007 log)

March Engineering (bonus SDT)

10-11 Mar 2007 Some more bonus SDT observing during engineering time. A few cirrus clouds; seeing never better than 1.2 arcsec.

11-12 Mar 2007 Some more bonus SDT observing during engineering time. Crystal clear; Seeing 1.5 arcsec in U early ....improving to 1 arcsec in B at the end of the night. U, R and V data on LeoII. Lost several hours to various telescope and instrument problems. Some data on Fan Quasar Point 2, but not a complete set. Sample images in moonlight of Fasano A1668.

March 12-20 SDT Run

12-13 Mar 2007 Clear, calm weather, but numerous failures, especially of the shutter, cost us the entire night. No PartnerObserving

13-14 Mar 2007 Camera shutter still malfunctioning, camera taken off telescope and into the lab. No PartnerObserving

14-15 Mar 2007 Good seeing, crazy night of malfunctions. Set the record for most IssueTraks (10) during a single night of observing. Very little useful observing despite pristine subarcsecond conditions.

15-16 Mar 2007 A Full Night of PartnerObserving [without problems], variable seeing, non-photometric, but we finally got to work.

16-17 Mar 2007 Spectacular seeing at the start, then on and off the rest of the night. Non-photometric.

17-18 Mar 2007 Heavy overcast and warm, break in clouds around 1:30am but poor (1.5-2" seeing), then 2:30am handed off for the Pluto Occultation

18-19 Mar 2007 Moving high thick clouds all night, a little observing in holes, only some of which was successful.

19-20 Mar 2007 Good seeing, an episode of great seeing, and then a lot of bad seeing. Worked all night, non-photometric due to high-thin cloud.

20-21 Mar 2007 Late start due to telescope work, then pretty good observing despite earlier cloud threat followed by rapid deterioration of seeing accompanied by clouds and high winds. Only a little observing was possible.

April Engineering (bonus SDT)

10-11 Apr 2007 Some more bonus SDT observing during engineering time. Clear, but seeing poor due to cold front. M81_BV data in 1.4 arcsec seeing.

April 11-23 SDT Run

11-12 Apr 2007 Some clouds and high wind at sunset. Observed Dwarf/mpobj3_1 and GRB070412, then clouds and wind took the rest of the night.

12-13 Apr 2007 Wind, Snow, and icy fog kept the telescope closed all night.

13-14 Apr 2007 It was snowing at 16:30. The clouds cleared up by midnight, but humidity kept us closed all night. We took some Dome flats.

14-15 Apr 2007 Sky is clear and humidity is nominal at 16:00. Could it be that we will finally take some data? Seeing variable from 0.6 to 1.5 arcsec. LBC shutter fails intermittently starting at 03:00 MST.

15-16 Apr 2007 Sky is clear, but the air is hazy and there's a lot of dust from the wind, so we're calling it non-photometric. The LBC shutter got stuck repeatedly, eventually we couldn't do anything about it.

16-17 Apr 2007 Cloudy skies at dusk, but cleared by 21:00. We are not open; they are attempting to fix the LBC shutter.

17-18 Apr 2007 Clear at sunset, but did not open until 04:00 MST because of work on the LBC shutter.

18-19 Apr 2007 Mostly clear, but high winds prevent opening until 20:00 MST. LBC shutter is working OK.......Finished Dwarf/mpobj_2. Seeing from 0.9 to 1.7 arcsec. Closed for high wind at 00:26 MST.

19-20 Apr 2007 Clear, worked all night with no LBC shutter issues, and no telescope issues. Seeing 1 arcsec until 2AM, then degrading 1.5-2 arcsec. Lost 60 min to misc LBC and Firefox problems.

20-21 Apr 2007 Clouds, Wind, Snow. Didn't open.

21-22 Apr 2007 Clear, worked all night. No problems at all until 3AM, then problems with Archive and telescope cost us 1.5 hours. Seeing was between 0.6 and 0.8 arcsec FWHM for several hours. Got GRB and M81 data early. Got good data on several Quasar fields and a bunch of Zaritsky photometric fields. Weaklensing data on A1835 probably won't be very useful because of variations in focus during the OB.

22-23 Apr 2007 Cloudy at the beginning, but clear after few hours. Worked all night. Seeing was between 0.9 and 1.3 arcsec. We got GRB070419a, Trilling fields and M81 in the first part of the night. Dwarf Can Ven in the R band, we complete Quasar Point 2, NGC6503.

23-24 Apr 2007 Snow storm. Then strong wind and clear. Probably photometric night. Seeing between 1.0 and 1.6 arcsec. Finish the Trilling field observations. Observed Quasar Point6. Then seeing worse to 1.5 arcsec and we have no more candidate in the SDT queue. Try NGC4826 and NGC5585. Photometric standard field and Sky flat at the end of the night.

May Engineering (bonus SDT)

09-10 May 2007 Some more bonus SDT observing during engineering time. Clear, with pretty good seeing. 5000 sec on GRB070419a with seeing from 0.8 to 0.9 arcsec. M81BV data. 2 OBs from Lockman Block 3, but seeing 1.3 - 1.4 arcsec as they set. NGC5474.

10-11 May 2007 Photometric. Lost first two hours to telescope problems. Finished Lockman Block 3 in marginal seeing. Lost one hour to LBC-Blue filter wheel problem. Good data set on IC4182 in 0.8 arcsec seeing. Then worked on pointing and collimation for a while.

11-12 May 2007 Broken clouds with poor seeing. Spent a couple hours on software issues. Worked on Lockman block 4. Finished Pizzella observations of NGC5719. Filled in with non-SDT field UGC10214. Zaritsky/NGC5585 with 1.2 - 1.5 arcsec images.

12-13 May 2007 We didn't open because of clouds until midnight. Observed Herc_R and some UrsaMinor dwarf fields in variable seeing.

14-15 May 2007 Checked focus offsets relative to V --- ran series of Vfastextra and finefocus sequences. Confirmed values for B, Uspec, U.

15-16 May 2007 Night started with thick clouds - seeing very poor until about 1-2 hrs before morning twilight. Attempted grb070419a (7-10pxl fwhm) and then did two OBs on RXJ1716... RXJ1716_CLASS_B_2.ob and RXJ1716_CLASS_B_3.ob with seeing about 1.3-1.5". Just before morning checked focus offsets for g and r relative to V and find these higher than the values entered in the configuration file. Have updated the configuration file with these new offsets: g-SLOAN 0.068 and r-SLOAN 0.07. Measured in ~1" seeing but through repeated alternating of fine focus sequences at V with those at r and g so as to have a frequent reference measurement, and the g and r values of ~0.07 repeated.

May 16-22 SDT Run

16-17 May 2007 Rain and snow or hail at the beginning of the night. Though humidity dropped to about 70-80% building was still dripping and there was indication of possible fog around still - did not open.

17-18 May 2007 Rain heading towards LBT + stars overhead ?!? initially. Then clearing, but telescope MCS control hung, and stuck for 2hr. Clouds again at midnight, with poor seeing. Then at 2am, reasonable clear, got 1hr on science before twilight,

18-19 May 2007 Some big clouds passing during twilight at the begining, then becoming clear. After a glitch with the LBC archive server, we got the entire night with seeing at ~1" (U). (lockman, GRBs, NGC6946, A2124)

19-20 May 2007 a very smooth night with no glitch & lots of data taken under seeing 1" or better- clear but not photometric. (M81, Lockman, NGC6946, RX1716, GRBs, Fan_qso) Listing of filename,OBname,object,filter,exptime,ra_off,dec_off

20-21 May 2007 Night could be used from dusk till dawn. Moon during first half. Seeing around 1.1" Clear and partly photometric (GRBs, M92, Lockman VLA, RXJ1716, NGC6946 )
Listing of filename,OBname,object,filter,exptime,ra_off,dec_off

21-22 May 2007 Moon even worse -- Seeing good at start but degraded, strong wind and high humidity after midnight. Some data taken, closed at 2AM.
Listing of filename,OBname,object,filter,exptime,ra_off,dec_off

22-23 May 2007 First quarter moon, but we think we can squeeze in one more observation of GRB070518. Closed by humidity.

June Engineering (bonus SDT)

08-09 Jun 2007 Got an M81 observation early. Spent rest of night in 0.6 to 0.7 arcsec seeing sorting out the filter focus offset issues. Adjusted the LBCFPIA alpha coefficient to make the Vfastextra focus consistent.

09-10 Jun 2007 Got an M81 observation early. Finished measuring all the filter focus offsets relative to Vfastextra. Took 20 x 200 sec of GRB070518 in 0.7 arcsec seeing while adjusting collimation to minimize the field-dependent astigmatism.

June 10-17 SDT Run

10-11 Jun 2007 Two Lockman fields with 1.5 arcsec images at 1.5 airmasses. CVen_B and CVen_V , CVen_2_B, UMi2_2_B, UMi2_B and UMi3_2_B with1" seeing

11-12 Jun 2007 Dome not opened due to bad weather (snow) in the evening, and high humidity during the second part of the night

12-13 Jun 2007 Clear Sky all night, seeing <1 arcsec all night, best seeing 0.7 arcsec,
Objects observed:
KochanekM81BV, KochanekM81V, N5024_shortUV, N5024_deepUV, M92shortR _F1, M92_deepR_F1, M92_deepB_F1, M92_deepU_F1, Umi3_V, Umi3_B, grb070518, hetej1900

13-14 Jun 2007 Clear Sky all night, seeing <1.5 arcsec all night, best seeing 1.0 arcsec,
Lost 1hour to instrument technical problems
Objects observed:
KochanekM81BV, KochanekM81V, RXJ1716_CLASS_B_2, RXJ1716_CLASS_B_3, NGC5907 in (r, g and U), M92_deepU_F2_1, M92_deepBR_F2_1 (partly)

14-15 Jun 2007 Thin clouds over the entire sky all night long, seeing <1.8 arcsec all night, best seeing 1.2 arcsec
Objects observed:
C2574_B1, Umi4_2_B , Umi4_B, Umi4_V, SN2006E1, SN2005cf_r1, SN2005cf_r2, RXJ1716_CLASS_B_8, M92_deepBR_F2_1, M92_short_F1, M92_short_F2, (M92_deepBR_F2_2)

15-16 Jun 2007 Thin clouds over the entire sky all night long, close dome before morning twighlight due to thick clouds , seeing <1.8 arcsec all night, best seeing 0.8 arcsec average 1.3 arcsec
Objects observed:
Umi5_2_B, Umi5_B, Umi5_V, (NGC5585 redo) , M92_deepBR_F2_2, M92_deepBR_F2_3, M92_deepBR_F2_4

16-17Jun 2007 Open for a while, but no serious data beyond art shots of NGC5466. Closed at midnight because of thickening clouds.

17-18 Jun 2007 Clear Sky all night, average seeing <1.5 arcsec all night, best seeing 1.0 arcsec, worst 2.5 arcsec,
Lost 1hour to instrument technical problems
Objects observed:
KochanekM81V (seeing 2.5), Umi6_2_B , Umi6_V, Umi6_B, Umicn1_B, Umicn1_V, Umicn1_2_B, M92_deepBR_F1, M92_deepU_F1, M92_deepU_F2_6, M92_deepBR_F2_6, grb070518(x2)

June Post-SDT PartnerObserving

18-19 Jun 2007 Photometric? Took several STDs throughout night, some for VTesta. Shutter tests for VTesta. Attempted QQQ1432 field, seeing too poor. Obtained art shot of M101 (2000s Uspec, 300s g, 200s r). Deep imaging of IRAC Calibration field (4500s Uspec). Test Ugr images (30s ea) on NGC7331. Data stopped arriving in the /Repository at 104755.

19-20 Jun 2007 Lost night to telescope and instrument problems. No science data.

20-21 Jun 2007 LBC Problems at midnight, so no science data.

21-22 Jun 2007 Clearing after clouds in afternoon. Lost 2 hours to PLC problems opening shutter doors. Got some data on M92_R through broken clouds. Took another 10 x 200 sec on GRB070518 during the only bad seeing of the night. Took a test exposure on M16, but couldn't focus there. Got SS433 r' data in 0.6 - 0.7 arcsec seeing. Took another 8 x 200 sec on GRB070518 in 0.8 arcsec seeing. Took 4-rotation g' flats. Had trouble with LSS at end.

22-23 Jun 2007 Thin clouds and smoke early, clearing later. We adjusted LBC-Blue hub down 1 mm for summer focussing. Spent a couple hours measuring field-astigmatism and focus effects. Did 2 hours of M92_deepU in 0.9 arcsec seeing after moonset.

23-24 Jun 2007 B Sky Flats; One last data point for M81V with 5000 counts background and 1.2 arcsec images ; Work on Red-Blue Co-Pointing; HETE J1900.1-2455 20 x 200 sec r' with 1.1-1.8 arcsec images at 1.9 airmasses.; One U and two R exposures on M92; Asteroid 17081 JT; R Sky Flats. Lost 10 minutes to instrument and 5 minutes to telescope.

September Engineering

15-16 Sep 2007 Clouds but were able to open around midnight. Took focus sequences at 0, 180 degrees and extrafocal pupil images at a range of elevations: 90-25 deg and back up for collimation lookup table.

16-17 Sep 2007 Clouds and rain at start, fog later on. 21-23 Sep were also foggy.

October Engineering

03-04 Oct 2007 High humidity kept us closed all night. Some dark tests with LBC-Blue shutter open. Issues with LBC-Blue Trackers bias voltage.hsel20071004: hsel20071004

04-05 Oct 2007 Engineering time commandeered from SDT. Open for an hour to adjust pointing zeropoint, then closed by humidity. We did get Uspec and V skyflats. Issues with Trackers CCD controllers continue (see Issues #838 and #839).hsel20071005: hsel20071005

05-06 Oct 2007 More engineering time commandeered from SDT. Open for an hour to get SkyFlats. Had trouble with mirror panics (Actuator 140) and TCS when we tried to do collimation. Closed for rising humidity. Fog and wind kept us closed for the rest of the night.hsel20071006: hsel20071006

October 06 - 17 SDT PartnerObserving

06-07 Oct 2007 Open for an hour, close for humidity, watch baseball for 40 minutes, open again for pointing / rotator center data. Start SDT observations under clear sky and dropping humidity at 04:10 UT. Close again for humidity at 05:48. Open again at 10:34 when humidity drops suddenly. Observed Ryan asteroid fields in variable seeing, and CL0016_C in poor seeing. Listing of images - columns are filename, OB name, object name, image type, exposure time, filter, xditheroffset: hsel20071007

07-08 Oct 2007 Beautiful clear skies with dew point as low as -31degC. Of course this means that we have problems with HBS pumps and lose 45 minutes at the beginning. Initial seeing was as good as 0.9 arcsec, but then degrades to 3 arcsec, and then slowly comes back down to 1 arcsec. Discovered a problem with the primary mirror oscillating. Worked on STDs fields and finished Ryan asteroid fields, then Andromeda XIV. Sub-arcsec images (per airmass) for the last couple hours. Jet crashed at 11:00, so we lost telescope tracking for 20 min. Listing of images - columns are filename, OB name, object name, image type, exposure time, filter, xditheroffset: hsel20071008.

08-09 Oct 2007 Sky is photometric, but the night starts off with 3 arcsec seeing. Observed AndXIV with 1.7 arcsec seeing. Lost 15 minutes to cranky actuator #140. Seeing continues worse than 2 arcsec, so we did large numbers of photometric standard fields, and a second pass over the Ryan asteroid fields. Seeing starts to improve around 3AM, so we got some AndXIV data in 1.7 arcsec seeing. MCSPU crashed again at 4AM costing us 15 minutes. By the end the seeing had improved to 1.8 arcsec at airmass 1.3, and some thin clouds had started to roll in. Listing of images - columns are filename, OB name, object name, image type, exposure time, filter, xditheroffset:hsel20071009.

09-10 Oct 2007 Rain at sunset. Problem with LBC archive filled up. Riccardo Smareglia fixed the problem and freed space. Some trouble with the mirror coming up only at 95%. Opened at 09:30 but after 40 minutes humidity raised again and we closed again.

10-11 Oct 2007 Night looks clear if not photometric. Few flats taken at dusk because there was still ongoing work in the dome. We had problems with telescope tracking and balancing at the beginning of the night. About 1:30 hr lost to fix the problem. Extra weights added to recover balancing. Then mirror problems required remote working from Tucson. Lost approx. other 50 min. At about 5UT moved to Andromeda XIV and completed the B,V program with seeing conditions ranging from 1.2-1.4 at the beginning to 0.9-1.0 at the end. Then tried to start Clementini program but seeing had already degraded to 1.3-1.4. Then we finished the night program with StanekM33, KochanekM81 BV and Nonino MS0451_C_2, with seeing between 1.2 and 1.7. After a standard field (SA95), we took some flat in V, B, Uspecial. Listing of images - columns are filename, OB name, object name, image type, exposure time, filter, xditheroffset:hsel20071011.

11-12 Oct 2007 Nice and clean sky. Problems at the beginning of the night to set the mirror. Actuator 140 fixed but gave problem so it was deactivated again. Seeing conditions good, always around 1.0-1.1 arcsec with peaks as good as 0.7 arcsec. Fields/programs observed: Schwope, Clementini (problem with tracking and guiding, so few data were secured), Zaritsky/Pasquali (the photometric OBs), Giallongo, Quirrenbach, Kochanek. Not many standards taken because at the end of night the azimut encoder needed to be reset so we have only two fields taken at the beginning and during the night. Flats taken in Uspec, U, B and V. As a notice to the next observers, we buiilt up an OB for the standard SA98, the richest in star density, with PA of -45 to have almost all the stars in the Landolt list in the field of view, with a decent even distribution all over the four chips. We wanted to test it at the end of night when we experienced the problem with the encoder. All in all, it was a candidate photometric night. Listing of images - columns are filename, OB name, object name, image type, exposure time, filter, xditheroffset:hsel20071012.

12-13 Oct 2007 Clear sky with some thin clouds at horizon. Problems with telescope encoders and brakes at the beginning of the night. Lost about 2 hrs. Usual difficuly in achieveing a good image analysis, we needed about 6 runs changing some cefficients manually before having a good result. Seeing not bad but encoders kept having problems. Some images elongated by about 24arcsec. Lost other 30 min fixing the problems, due to azimuth encoders. Seeing went bad since, about 1.4-1.7. Programs executed: Schwope (aborted for seeing and bad tracking), Majewski, Clementini (interrupted for bad seeing), Sand (seeing quite poor). Some standards at the end although the night has not been photometric. Listing of images - columns are filename, OB name, object name, image type, exposure time, filter, xditheroffset: hsel20071013.

13-14 Oct 2007 Some clouds at sunset. Humidity around 80 %. Major problems with the telescope, about 2.5 hrs lost. DIfficult to achieve a good image analysis at the beginning of the night, it took more than 1 hr and a call to John Hill. Seeing average, around 1.1-1.5, better at the end of night. Executed OBs for Schwope, Clementini, Stanek, Sand. One standard field at the end of the night to monitor the zero points. Listing of images - columns are filename, OB name, object name, image type, exposure time, filter, xditheroffset:hsel20071014.

14-15 Oct 2007 Clear night with 1-1.5 arcsec seeing. Spent 1+ hr at beginning of the night establishing focus. Finished 15 deg Ryan field with 3 epochs. Several Clementini/M31 OBs. One Schwope OB that needs to be checked by PI before being checked off. One M33 Stanek OB that inexplicably had very bad distortion across the field. Finished And_IX_B OB. One V-band M81 OB. And finally two Ursa Major II OBs. Good night with no real problems except for focus at beginning of the night. Fingers crossed for tomorrow... Listing of images - columns are filename, OB name, object name, image type, exposure time, filter, xditheroffset:hsel20071015.

15-16 Oct 2007 Clear night. Began with Schwope fields and then M31 Clementini fields until JET crash ends night around 1000 UT. Listing of images - columns are filename, OB name, object name, image type, exposure time, filter, xditheroffset:hsel20071016.

16-17 Oct 2007 Cruddy night. Did not open until ~2am due to humidity. Then seeing was 2+ arcsec. Oh well. Got some Ryan OBs done, but telescope problems kept us from getting three epochs. Bust. Listing of images - columns are filename, OB name, object name, image type, exposure time, filter, xditheroffset:hsel20071017.

17-18 Oct 2007 Clear, but seeing above 2 arcsec for nearly all the night. When seeing would dip, we would try Clementini or other programs, but the seeing would quickly return to ~2 arcsec. Did several Ryan sequences. Good bye SDT. Listing of images - columns are filename, OB name, object name, image type, exposure time, filter, xditheroffset, yditheroffset:hsel20071018.

November Engineering

[[05Nov2007][05-06 Nov 2007] A few images of Comet Holmes and M78 during technical observing........not photometric
I Attachment Action Size Date Who Comment
20070116.hselhsel 20070116.hsel manage 13 K 20 Feb 2007 - 23:48 OlgaKuhn 15/16 January 2007 log
20070117.hselhsel 20070117.hsel manage 6 K 20 Feb 2007 - 23:50 OlgaKuhn 16/17 January 2007 log
20070118.hselhsel 20070118.hsel manage 14 K 20 Feb 2007 - 23:50 OlgaKuhn 17/18 January 2007 log
20070214.hselhsel 20070214.hsel manage 6 K 20 Feb 2007 - 23:51 OlgaKuhn 13/14 February 2007 log
20070216.hselhsel 20070216.hsel manage 22 K 20 Feb 2007 - 23:54 OlgaKuhn 15/16 February 2007 log
20070217.hselhsel 20070217.hsel manage 14 K 20 Feb 2007 - 23:51 OlgaKuhn 16/17 February 2007 log
20070218.hselhsel 20070218.hsel manage 8 K 21 Feb 2007 - 00:01 OlgaKuhn 17/18 February 2007 log
20070219.hselhsel 20070219.hsel manage 7 K 21 Feb 2007 - 00:01 OlgaKuhn 18/19 February 2007 log
20070221.hselhsel 20070221.hsel manage 22 K 21 Feb 2007 - 13:43 OlgaKuhn 20/21 Feb 2007 log
20070222.hselhsel 20070222.hsel manage 17 K 22 Feb 2007 - 12:00 OlgaKuhn 21/22 Feb 2007 log
20070223.hselhsel 20070223.hsel manage 14 K 23 Feb 2007 - 12:47 OlgaKuhn 22/23 Feb 2007 log
20070225.hselhsel 20070225.hsel manage 20 K 25 Feb 2007 - 20:46 OlgaKuhn 24/25 Feb 2007 log
20070226.hselhsel 20070226.hsel manage 7 K 27 Feb 2007 - 02:13 OlgaKuhn 25/26 Feb 2007 log
20070227.hselhsel 20070227.hsel manage 8 K 27 Feb 2007 - 18:26 OlgaKuhn 26/27 Feb 2007 log
hsel0522EXT hsel0522 manage 10 K 22 May 2007 - 19:26 OlgaKuhn  
hsel20070520EXT hsel20070520 manage 9 K 21 May 2007 - 16:54 OlgaKuhn  
hsel20070521EXT hsel20070521 manage 13 K 21 May 2007 - 16:52 OlgaKuhn  
hsel20071004EXT hsel20071004 manage 2 K 17 Oct 2007 - 17:52 OlgaKuhn  
hsel20071005EXT hsel20071005 manage 9 K 17 Oct 2007 - 17:53 OlgaKuhn  
hsel20071006EXT hsel20071006 manage 16 K 17 Oct 2007 - 17:54 OlgaKuhn  
hsel20071007EXT hsel20071007 manage 17 K 17 Oct 2007 - 17:55 OlgaKuhn  
hsel20071008EXT hsel20071008 manage 35 K 17 Oct 2007 - 17:59 OlgaKuhn  
hsel20071009EXT hsel20071009 manage 38 K 17 Oct 2007 - 17:54 OlgaKuhn  
hsel20071011EXT hsel20071011 manage 16 K 17 Oct 2007 - 17:55 OlgaKuhn  
hsel20071012EXT hsel20071012 manage 39 K 17 Oct 2007 - 17:56 OlgaKuhn  
hsel20071013EXT hsel20071013 manage 12 K 17 Oct 2007 - 17:56 OlgaKuhn  
hsel20071014EXT hsel20071014 manage 15 K 17 Oct 2007 - 17:56 OlgaKuhn  
hsel20071015EXT hsel20071015 manage 21 K 17 Oct 2007 - 17:57 OlgaKuhn  
hsel20071016EXT hsel20071016 manage 12 K 17 Oct 2007 - 17:57 OlgaKuhn  
hsel20071017EXT hsel20071017 manage 11 K 17 Oct 2007 - 17:57 OlgaKuhn  
hsel20071018EXT hsel20071018 manage 17 K 18 Oct 2007 - 20:23 OlgaKuhn  
Topic revision: r167 - 15 Oct 2009, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback